The ratio between the sum of 'n' terms of two arithmetical progressions is (7n+1):(4n+27). Find the ratio of their eleventh term.
Ans: 148/111

Answer by t0hierry(194)

n=21
7n+1 = 148
4n+27= 84+27=111